# Bias Is a Subspace, Not a Coordinate: A Geometric Rethinking of Post‑hoc Debiasing in Vision-Language Models

### v1

@@ 5.2.3 Text-to-Image Generation
Given a text prompt, the model generates an image. We study two failure modes: (i) producing the wrong gender when the prompt specifies one (e.g., generating a woman for “a photo of a man who works as a/an [profession]”), and (ii) producing a gender-skewed distribution for neutral prompts (e.g., always generating men for “a photo of a person who works as a/an [profession]”). We use profession-based prompts spanning 83 occupations derived from Bias-in-Bios. For each prompt, we generate 10 images with different random seeds.
@@ 6.1 Effectiveness of SPD on SFID’s Limitations
Table 3 summarizes the results of the evaluation procedure described in Sec. 5.1. For the proposed SPD method, when the number of removed subspace directions is r=1r{=}1, the accuracy of the target attribute remains almost unchanged, indicating that attribute information is not concentrated in a single direction but distributed across multiple discriminative directions. Increasing $r$ to five substantially reduces the probe accuracy for the target attribute while the non-target attributes change by less than one percentage point in most cases. With a moderate projection depth, SPD effectively removes target-attribute information while introducing minimal collateral impact. This indicates that SPD achieves more complete debiasing while substantially alleviating the feature entanglement problem inherent in coordinate-level methods. As $r$ increases to 10, the target accuracies decrease further, but the non-target accuracies also begin to drop more noticeably. This pattern suggests that the lower-ranked directions are more entangled with non-target attributes and semantic signals, so deeper projection starts to remove task-relevant information.
@@ 6.2 Better Performance Across Downstream Tasks
Zero-shot Classification task (Tab. 4): On this task, SPD consistently achieves the lowest demographic-parity gaps across all backbones with only minor accuracy movement. Removing the subspace, rather than isolated coordinates, reduces disparities more reliably than other baselines, yielding the best fairness-utility trade-off overall. Taken together, these results show that SPD achieves the best fairness–utility trade-off among all compared approaches.
@@
Across diverse datasets and downstream tasks, SPD achieves the strongest bias mitigation while preserving semantic consistency. These results demonstrate that the learned bias subspace exhibits stable and transferable invariance across distributions, enabling SPD to generalize robustly across modalities and tasks, outperforming coordinate-level post-hoc methods.
6.3 Ablations Study
@@ 7 Conclusion
We revisit post-hoc debiasing for VLMs and reveal three key systematic deficiencies in coordinate-level editing approaches, including residual bias leakage, semantic entanglement, and poor cross-distribution generalization. We further find that demographic bias in VLM embeddings is fundamentally subspace-structured rather than coordinate-sparse. Building on this geometric insight, we propose SPD, a theoretically grounded and training-free framework that removes all linearly decodable attribute components and reinjects a neutral on-manifold representation to maintain semantic fidelity. Across diverse downstream tasks, SPD consistently achieves substantial reductions in demographic disparity while preserving task utility. These results confirm the subspace-structured nature of bias and demonstrate that exploiting this geometry enables robust and effective debiasing across distributions and modalities.

### v2

@@ 5.2.3 Text-to-Image Generation
Given a prompt, the model generates an image. We study two failure modes: (i) producing the wrong gender when the prompt specifies one (e.g., generating a woman for “a photo of a man who works as a/an [profession]”), and (ii) producing a gender-skewed distribution for neutral prompts (e.g., always generating men for “a photo of a person who works as a/an [profession]”). We use profession-based prompts covering 83 occupations from Bias-in-Bios. For each prompt, we generate 10 images with different seeds.
@@ 6.1 Effectiveness of SPD on SFID’s Limitations
Table 3 summarizes the results of the evaluation procedure described in Sec. 5.1. For the proposed SPD method, when the number of removed subspace directions is r=1r{=}1, the accuracy of the target attribute remains almost unchanged, indicating that attribute information is not concentrated in a single direction but distributed across multiple discriminative directions. Increasing $r$ to five substantially reduces the probe accuracy for the target attribute while the non-target attributes change by $1\%$ in most cases. With a moderate projection depth, SPD effectively removes target-attribute information while introducing minimal collateral impact. This indicates that SPD achieves more complete debiasing while substantially alleviating the feature entanglement problem inherent in coordinate-level methods. As $r$ increases to 10, the target accuracies decrease further, but the non-target accuracies also decrease noticeably. This suggests that the lower-ranked directions are more entangled with non-target attributes and semantic signals, where deeper projection removes task-relevant information.
@@ 6.2 Better Performance Across Downstream Tasks
Zero-shot Classification task (Tab. 4): On this task, SPD consistently achieves the lowest demographic-parity gaps across all backbones with only minor accuracy movement. Removing the subspace, rather than isolated coordinates, reduces disparities more reliably than other baselines, yielding the best fairness-utility trade-off overall. Overall, SPD provides consistent fairness gains and improves the fairness–utility Pareto frontier relative to all baselines.
@@
Across diverse datasets and downstream tasks, SPD achieves the strongest bias mitigation while preserving semantic consistency. These results suggest that the learned bias subspace exhibits more stable cross-distribution behavior, partially mitigating Assumption A2 and enabling SPD to generalize more robustly across modalities and tasks compared to coordinate-level post-hoc methods.
6.3 Ablation Study
@@ 7 Conclusion
We revisit post-hoc debiasing for VLMs and reveal three key systematic flaws in coordinate-level editing approaches, including residual bias leakage, semantic entanglement, and poor cross-distribution generalization. We further find that demographic bias in VLM embeddings is fundamentally subspace-structured rather than coordinate-sparse. Building on this geometric insight, we propose SPD, a theoretically grounded and training-free framework that removes controllable attribute-predictive directions and reinjects a neutral mean component. Because protected attributes can correlate with semantic factors, projection depth $r$ governs a fairness-utility tradeoff; across the evaluated datasets and tasks, SPD reduces demographic disparity metrics while maintaining competitive task performance.
Acknowledgement
We thank the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University for computational support.